Raith Status Friday Septrember 17, 2004 1330 hours.

James Conway jwc at snf.stanford.edu
Fri Sep 17 14:15:13 PDT 2004


Greetings Raith Community:

I had a large amount of time on the system this week and was able to 
complete two difficult pattern writes over the last two days. These are 
the wafers that I had been trying to write for nearly a month. A number 
of simple test and check outs were performed in parallel to this work.

The RAITH 150 system is working well and appears to be within 
specification for stitching and better than specification for overlay. I 
will be repeating the PCM performance testing we completed August 28, 
2004 to confirm we are still were we were then. As this work is 
completed I will report the results and post images on the bulletin 
board superseding the work done in May 2004.

LEO /Zeiss SMT will return early next week for adjustments to and 
checking the settings on the E/O and column calibrations we performed 
two weeks ago.  We will also attempt to bring down the Zoom V value to 
levels farther from the Zoom V Limit value. I have been able to complete 
a number (9X3) write field alignments without encountering the Sweep Sum 
exceeds limits error.  Several users whom encountered problems during 
write field alignment had not measured and adjusted the WD to 5 mm using 
the SEM to measure WD. IN one case they were only at 3 mm!

The FE-Gun tip has run in nicely and Gun vacuum EHT ON is in the 8E-010 
T region.
The system has been tested this week to 25 kV  with extended writes 
performed at 20 and 10 keV.  I will perform some pattern writes next 
week at 20,25, and 30 kV to ensure beyond any doubt that we can start 
general operations at this acceleration voltage for any users wishing to 
do so.

I recalibrated the Piezo height controls using a bare Si wafer on the 
ElectroStatic Chuck (ESC) and the new calibration for centered post is 
within 0.149 mm across a six inch wafer.  I will be making mechanical 
adjustments to the ESC and will recalibrate the piezo's once again.  At 
that point we should be able to level samples on the ESC and not exceed 
the range of the piezo's travel in Z.  Following that I will evaluate 
the leveling of the Standard Sample holder as it sits onto the kinetic 
mounts of the Piezo post and see if there are actually any issues with 
this holder being level or not.

Made measurements of the Beam Current at 10 keV and 5 mm Working 
distance for all the apertures.
We are in the middle of the decent of Beam Current as we close these 
apertures with hydrocarbon contamination.  Overall we are doing well on 
this issue and beam speeds are near optimum for all apertures on 100 nm 
thick 2% 950K MW PMMA; allowing us to use suitably small area steps and 
still have Beam speed below 10 mm/sec.

Beam Current (Beam I) measurements I made today:

APT            Beam I Today             Beam I after May PM and aperture 
plate replacement            Delta
(um)              ( nA )                           (nA  aperture values 
just installed)                                          (% of Beam I new)

120               2.345                              4.842               
                                                                  48
60                 0.6182                           1.025               
                                                                   60   
                                   
30                 0.1339                           0.233               
                                                                   57
20                 0.0699                           0.085   6-15-2004 
first value recorded in logbook?             82
10                 0.0168                           0.0459            
                                                                    36  
                                     7.5                 0.0078         
                  0.011                                                
                                 70

Conclusions:  We are approximately 50 % into our PM interval with 
acceptable beam currents for all apertures albeit slightly on the low 
side for the 10 um aperture.  This is consistent with the usage we have 
on the system for each aperture, particularly considering the large 
number of EBL writes that have been performed this summer.
  The 10 um aperture beam I is reducing more rapidly than during the 
previous PM interval last year.  This value may be indicative of a 
greater contribution of contamination from samples by users working with 
this aperture than the 30 um aperture.

Thanks to the many users which have been diligently recording their Beam 
Currents in the logbook.  This information is very useful to us and 
helps me greatly to monitor the condition of the system.

System is up and in Green Light for your usage.  If anyone is having any 
problems whatsoever on the system please see me during my office hour.

Thank you for your support and have a pleasant weekend.

James Conway

Upcoming in October:
  RAITH Group XI a four day intensive short course will be conducted the 
week of October 12 - 15, 2004
  There are two slots remaining in this class.  Please maximize your 
utilization of the system coming into this period.






-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/raith/attachments/20040917/9f715615/attachment.html>


More information about the raith mailing list